# Computer-Controlled 3D Freeform Surface Weaving

### v1

@@ 1 Introduction
The history of textile fabrication is probably as old as human historyadanur2020handbook . With the development of civilization, the function of textile products has been extended from keeping warmth to smart wearable devices in applications such as aerospace engineering kelkar2006structural and biomedical engineering di2006magic ; 9852702 . There are growing demands for 3D shaping new composite materials mouritz1999review and e-textile komolafe2021textile ; levitt20203d , where the geometry of fabrics could possibly be formed by four different methods including knitting mccann2016compiler ; narayanan2018automatic ; liu2019computational ; igarashi2008knitting , weaving bilisik2010multiaxis ; wu2020automatic ; ren20213d , feltingchen2015advances and robotic deposition / winding  shirinzadeh2004fabrication ; shirinzadeh2007trajectory ; SORRENTINO201743 .
In the fabrics formed by felting, threads / fibres are randomly distributed (see the right of Fig.1) – i.e., not controllable. The approaches of robotic direct deposition can only fabricate structures with weakly bonded threads / fibres – i.e., the threads are ‘stitched together’ by chemical but not mechanical bonding in a manner similar to filament based 3D printing shirinzadeh2004fabrication ; shirinzadeh2007trajectory , and the winding technique is mainly limited to the surfaces obtained by revolution SORRENTINO201743 ; li2021Winding . To overcome these problems, a new technology needs to be developed for fabricating woven structures with controllable locations of threads on 3D surfaces. Note that, for the sake of explanation, we simply call each unit structure of woven fabrics as a stitch in the rest of this paper. Fig.1 shows the stitch structures in weaving (left), in knitting (middle), and the unstructured threads in felting (right). Generally, all structured stitches need to be mapped into the elements of a two-dimensional grid layout to enable the manufacturing process. Such a map is named as weaving map (see Fig.2(b) for an example). Weaving process results in interlaced warp and weft threads that are nearly perpendicular to each other adanur2020handbook – fabrics can therefore be formed by threads with high bending-stiffness. This enables many applications using threads with advanced physical functions such as carbon fibres, optical fibres, and conductive wires (ref. mouritz1999review ; komolafe2021textile ; levitt20203d ).
@@
Three-dimensional freeform shapes can be formed in knitting by the operations to increase or decrease the number of stitches, which has been commonly adopted in hand knitting for simple shapes such as a rounded cap since the 16th century spencer2001knitting . Different from weaving, the inter-looped stitches in knitting are formed by warping a continuous filling thread throughout the whole fabric-forming process. More complicated geometry was considered in computer graphics by employing increase-decrease shaping for hand knitting igarashi2008knitting or short-row shaping for machine knitting liu2021knitting . Increase / decrease shaping changes the local width of the fabric by adding an extra loop in a row of stitches (or trapping a loop to more than one loop in the previous row). Short-row shaping changes the local height of the fabric by knitting partial rows in regions needing more stitches. Modern computational techniques have been developed to generate knitting code for realizing desired shapes on knitting machines, including the compiler mccann2016compiler , the visual programming tool narayanan2019visual , and the algorithm for controlling the distribution of elasticity liu2021knitting . However, highly curved loops formed in knitting make it challenging to use materials with high bending-stiffness (see the middle of Fig.1). An alternative manufacturing method is needed.
We aim at developing a new technology for computer-controlled fabrication of 3D freeform shapes by woven structures. Different from the existing works on weaving 3D solid geometry like height-fields (ref. wu2020Woven3D ), the weaving technique developed in our work focuses on producing 3D surface geometry. 3D surface weaving products can be widely seen in our daily life (e.g., woven baskets and woven chairs), which however are mainly hand-made. Ren et al.ren20213d presented an optimization-based approach to design woven structures by an algorithm to compute the ribbon’s planar geometry. However, the manufacturing process has not been considered. The challenges for realizing a computer-controlled 3D surface weaving process include:
@@
To the best of our knowledge, these do not exist in literature although the demanded algorithm shares some similarity to our earlier work of knitting map generation (ref. liu2021knitting ).
@@ 2 Weaving 3D Freeform Surface
Jacquard devicemiura1964jacquard on a flat weaving machine, which is originally used to product flat embossed patterns on fabrics, is first modified to enable the function of partial weaving by controlling different up and down states of warp threads (i.e., those along the column directions). As a result, different numbers of warp threads can be woven by the weft thread (i.e., along the row direction) starting from / ending at different locations – see Fig.2(b, c) for an example and Sec. 3.1 for more details. Secondly, considering that the shape of 3D woven fabric is determined according to the lengths of the warp threads in different columns, different warp threads need to be controlled to ‘tighten’ or ‘release’ individually (Sec. 3.2). A new mechanism as a matrix of warp beams is developed to realize this function (see the right of Fig.3). How to effectively form the desired pulling operation while keeping the warp threads tight all the time during the weaving process is one of the most challenging tasks. A weaving mechanism is designed for this purpose (Sec. 3.3). The other challenging task is weft length control, which is realized in our system by dual robotic-arms equipped with cameras (Sec. 3.4).
Figure 3: Our computer-controlled 3D weaving system can fabricate woven structures to form a target 3D surface shape and it mainly consists of four hardware parts: 1) jacquard device, 2) warp beams, 3) weaving mechanism, and 4) shuttle moving mechanism. The weaving process can be completed automatically with the help of two robotic arms equipped with cameras. The weaving result as a hemisphere with embedded optical fibres is shown in the right-upper corner.
In the software aspect, an algorithm is developed to generate weaving maps for controlling the operations of our weaving machine (see Fig.2(d) for an example). The algorithm is based on first generating a knittable stitch mesh by using the geodesic field based method in liu2021knitting . The stitch mesh for knitting is converted into a weaving map by considering the jacquard device and the weaving mechanism employed in our system. Each weaving map is later converted into a set of operations defined as W-code to govern the machine weaving. Details of the algorithm can be found in Sec. 4.2.
@@ 3.1 Partial weaving by Jacquard device
Figure 5: Each warp beam is compounded of a frame a⃝, a coil of warp thread b⃝, a motor c⃝, and a pair of driving gears d⃝ and e⃝.
@@ 3.3 Weaving mechanism
Figure 6: A novel weaving mechanism (a) is invented in our system to form required ‘push’ of weaved stitches on the fabric by operating 15 components together. Specifically, the front piece 1⃝2 and the back piece 1⃝4 of the reed are fixed while the middle piece 1⃝3 can shift a bit along the weft direction driven by a pair of eccentric wheels 1⃝5. The detailed structure of the reed is given in (b).
@@ 3.4 Weft length control with the help of robotic manipulation
In our implementation, we mount a camera on the top of fabric clamp to capture the position of the weft thread using the edge extraction based algorithm – see the window views at the upper-left corner of Fig.8(c) and (d). Two TM900 robotic arms are used to drive the slender hooks as their end effector, where the camera mounted on a robotic arm can help to successfully capture the weft thread by the hook. This robot-assisted tightening operation can help form stitches with better dimension control on the resultant fabrics.
@@ 3.5 Control system
Figure 10: Given a target surface, we first generate a geodesic distance-field by using the central line of the piece as source (a). A stitch mesh for short-row shaping based knitting as shown in (b) can be generated from the isocurves of the geodesic distance-field by optimized quadrangles / triangles (details can be found in liu2021knitting ). The stitch mesh can directly map into (c) a knitting map $K(\cdot,\cdot)$, which is converted into (d) a weaving map $W(\cdot,\cdot)$ and the machine code with its structure / syntax as explained in (e) to supervise the operations of our weaving machine.
@@ 4.1 Knitting map generation
The stitch mesh generated by the algorithm in liu2021knitting can satisfy all these requirements. As a result, every stitch can be mapped to a location in the 2D knitting map by its row index and its location in the row – see Fig.10(b, c) for an example.
@@ 4.2 Converting knitting map into weaving code
Figure 11: We tested our 3D surface weaving machine on four different examples, where their weaving maps are given in (a). The results of physical fabrication are shown in (b). Our weaving system enables the flow of electric current over the freeform surfaces of fabrics by using conductive threads in (c).
@@
Trgl. Thread ${}^{{\dagger}}$ Comp. ${}^{{\ddagger}}$ (sec.)
@@
${}^{\dagger}$ Tests with different number of threads were conducted to demonstrate the scalability of our computational method although the maximally allowed number of warp threads on our prototype machine is only 100.
${}^{\ddagger}$ The computing time of W-code generation includes both the step of weaving map generation and the step for converting it into W-code.
@@ 5.1 Statistics of computation and fabrication
The computational statistics of these examples have been listed in Table 1, where the examples are tested by using threads in two different resolutions. As it can be observed, the computation on all tests can be completed within 10 seconds and the major bottleneck is the step of knitting map generation liu2021knitting . The complexity of the knitting map generation algorithm is $O(AB\log B)$ with $A$ and $B$ being the maximal number of stitches along the row and the column directions on the knitting map respectively.
@@ 5.3 Analysis of geometric errors: 3D shape
One of the most important purposes of this work is to reach a high level of geometric accuracy on the 3D surface shape of the produced woven fabrics. In the conventional textile industry, a flattening algorithm (e.g., TASE2011 ; CAD2005 ) was employed to compute a (trimmed) 2D panel shape to fit onto the desired 3D object (e.g., human bodies in the apparel industry). However, as 3D surfaces are in general not flattenable (or named as developable in geometric modeling), wrinkles will appear at regions where the curvature is large and leftovers are inevitable. For a better fit, smaller pieces of the flat-woven fabric are generated and stitched together, which involved time-consuming sewing processes that are labor-intensive (e.g., zhao2023evolutionary ; zhang2019FabFormwork ; TVCG2008 ).
We conducted experimental tests to study the shape approximation errors on a variety of models. Both the skinned shape by flat woven fabrics and the shape from 3D surface weaving are compared with the target shape (as the ground truth). A structured light based 3D scanner sacnner is employed to generate the point clouds for each surface, which is also placed into a well-registered pose w.r.t. the ground truth shape. The shortest distances between the sampled points on the given surface to the scanned point clouds are employed as shape approximation errors, which are visualized as color maps and histograms in Fig.14. From the results of error analysis, we can easily conclude that models fabricated by our 3D surface weaving have much smaller errors in all examples – i.e., mainly less than 1mm as can be observed from the histograms. When the surface to be fabricated becomes more and more complex, the geometric error on fitting results by flat woven fabrics will increase significantly (see the triple-peak model in Fig.14 for an example).
@@ 5.6 Continuity of Threads
Our 3D surface weaving machine can help produce functional fabrics with many interesting applications in mechatronics. When conductive threads are used (as already demonstrated in Fig.11(c)), our 3D surface weaving approach enables the flow of electric current over the surfaces of fabrics so that they can work like soft PCBs. This has many potential applications such as functional garments 9788480 , motion sensing 9263352 , and ‘skin’ of cobot Pang2021CoboSkin . To enable this function, continuity on both the warp and the weft threads is required. While forming a 3D freeform shape by our surface weaving technique, we use a single thread for weft and multiple continuous threads for warp directions. Continuity is preserved on the resultant woven fabrics.
@@ 5.7 Limitations
One major limitation of our current implementation is that two robotic arms are employed for realizing the length control of weft threads. This significantly increases the cost and the complexity of the manufacturing system. A possible future work is to design a special mechanism equipped with force sensors to complete the tightening step for weft threads.
@@
Figure 19: An example of vest-front fabric made by using 1) bare stainless steel wire for warp threads and 2) silicone insulated copper wire for weft threads (left), which can be developed into a touch sensing interface by capacitance Lotters99capacitance . (Middle) the distribution of touch that is sensed by the grid of conductive threads, and (right) the corresponding touch visualized on the 3D surface.
In the examples listed above, the threads with high stiffness (such as optical fibres) are only employed as warp threads. This is because the weft thread needs to form U-turn at the end of each row. However, we can employ the in-lay technique to ‘insert’ an additional thread (with high bending-stiffness) between those neighboring rows formed by a continuous weft thread. An example can be found in Fig.18. Note that the in-lay fabrication needs manual operations on our prototype machine in our current experiment. However, when the U-turn at the end of each row is tolerable on materials with high bending-stiffness (e.g., metal wires), our method can produce fabrics by using these stiff materials in both warp and weft threads – see Fig.19 for an example by using these conductive grids to form a capacitance-based touch sensing interface Lotters99capacitance . Lastly, the abrasion of warp threads caused by the reed’s motion requires the threads to be waxed to prevent damage to its ‘core’.
There are also some limitation on the software aspect. Our current method of weaving-map generation heavily relies on selecting a good curve as source to generate geodesic distance-field. Different maps including failure cases can be generated when using different sources. One possible future research is to automate the source selection while preventing the failure cases and reducing the distortion of stitch meshes constructed from the geodesic distance-field.

### v2

@@ 1 Introduction
The history of textile fabrication is probably as old as human history[1]. With the development of civilization, the function of textile products has been extended from keeping warmth to smart wearable devices in applications such as aerospace engineering [2] and biomedical engineering [3, 4]. There are growing demands for 3D shaping new composite materials [5] and e-textile [6, 7], where the geometry of fabrics could possibly be formed by four different methods including knitting [8, 9, 10, 11], weaving [12, 13, 14], felting[15] and robotic deposition / winding  [16, 17, 18].
In the fabrics formed by felting, threads / fibres are randomly distributed (see the right of Fig.1) – i.e., not controllable. The approaches of robotic direct deposition can only fabricate structures with weakly bonded threads / fibres – i.e., the threads are ‘stitched together’ by chemical but not mechanical bonding in a manner similar to filament based 3D printing [16, 17], and the winding technique is mainly limited to the surfaces obtained by revolution [18, 19]. To overcome these problems, a new technology needs to be developed for fabricating woven structures with controllable locations of threads on 3D surfaces. Note that, for the sake of explanation, we simply call each unit structure of woven fabrics as a stitch in the rest of this paper. Fig.1 shows the stitch structures in weaving (left), in knitting (middle), and the unstructured threads in felting (right). Generally, all structured stitches need to be mapped into the elements of a two-dimensional grid layout to enable the manufacturing process. Such a map is named as weaving map (see Fig.2(b) for an example). Weaving process results in interlaced warp and weft threads that are nearly perpendicular to each other [1] – fabrics can therefore be formed by threads with high bending-stiffness. This enables many applications using threads with advanced physical functions such as carbon fibres, optical fibres, and conductive wires (ref. [5, 6, 7]).
@@
Three-dimensional freeform shapes can be formed in knitting by the operations to increase or decrease the number of stitches, which has been commonly adopted in hand knitting for simple shapes such as a rounded cap since the 16th century [20]. Different from weaving, the inter-looped stitches in knitting are formed by warping a continuous filling thread throughout the whole fabric-forming process. More complicated geometry was considered in computer graphics by employing increase-decrease shaping for hand knitting [11] or short-row shaping for machine knitting [21]. Increase / decrease shaping changes the local width of the fabric by adding an extra loop in a row of stitches (or trapping a loop to more than one loop in the previous row). Short-row shaping changes the local height of the fabric by knitting partial rows in regions needing more stitches. Modern computational techniques have been developed to generate knitting code for realizing desired shapes on knitting machines, including the compiler [8], the visual programming tool [22], and the algorithm for controlling the distribution of elasticity [21]. However, highly curved loops formed in knitting make it challenging to use materials with high bending-stiffness (see the middle of Fig.1). An alternative manufacturing method is needed.
We aim at developing a new technology for computer-controlled fabrication of 3D freeform shapes by woven structures. Different from the existing works on weaving 3D solid geometry like height-fields (ref. [23]), the weaving technique developed in our work focuses on producing 3D surface geometry. 3D surface weaving products can be widely seen in our daily life (e.g., woven baskets and woven chairs), which however are mainly hand-made. Ren et al.[14] presented an optimization-based approach to design woven structures by an algorithm to compute the ribbon’s planar geometry. However, the manufacturing process has not been considered. The challenges for realizing a computer-controlled 3D surface weaving process include:
@@
To the best of our knowledge, these do not exist in literature although the demanded algorithm shares some similarity to our earlier work of knitting map generation (ref. [21]). Considering the requirements that 1) the first row and the last row of weaving are located on the boundary of 3D surface patches and 2) only one weft thread is allowed to travel during the process, the 3D surface patch utilized in this paper should be an oriented 2-manifold with disk-like topology.
@@ 2 Weaving 3D Freeform Surface
Jacquard device[24] on a flat weaving machine, which is originally used to product flat embossed patterns on fabrics, is first modified to enable the function of partial weaving by controlling different up and down states of warp threads (i.e., those along the column directions). As a result, different numbers of warp threads can be woven by the weft thread (i.e., along the row direction) starting from / ending at different locations – see Fig.2(b, c) for an example and Sec. 3.1 for more details. Secondly, considering that the shape of 3D woven fabric is determined according to the lengths of the warp threads in different columns, different warp threads need to be controlled to ‘tighten’ or ‘release’ individually (Sec. 3.2). A new mechanism as a matrix of warp beams is developed to realize this function (see the right of Fig.3). How to effectively form the desired pulling operation while keeping the warp threads tight all the time during the weaving process is one of the most challenging tasks. A weaving mechanism is designed for this purpose (Sec. 3.3). The other challenging task is weft length control, which is realized in our system by dual robotic-arms equipped with camera together with a camera mounted above the weaving region (Sec. 3.4).
Figure 3: Our computer-controlled 3D weaving system can fabricate woven structures to form a target 3D surface shape and it mainly consists of four hardware parts: 1) jacquard device, 2) warp beams, 3) weaving mechanism, and 4) shuttle moving mechanism. The weaving process can be completed automatically with the help of two robotic arms equipped with camera together with a camera mounted above the weaving region. The weaving result as a hemisphere with embedded optical fibres is shown in the right-upper corner.
In the software aspect, an algorithm is developed to generate weaving maps for controlling the operations of our weaving machine (see Fig.2(d) for an example). The algorithm is based on first generating a knittable stitch mesh by using the geodesic field based method in [21]. The stitch mesh for knitting is converted into a weaving map by considering the jacquard device and the weaving mechanism employed in our system. Each weaving map is later converted into a set of operations defined as W-code to govern the machine weaving. Details of the algorithm can be found in Sec. 4.2.
@@ 3.1 Partial weaving by Jacquard device
Figure 5: Each warp beam is compounded of a frame ⓐ, a coil of warp thread ⓑ, a motor ⓒ, and a pair of driving gears ⓓ and ⓔ.
@@ 3.3 Weaving mechanism
[…]
@@ 3.4 Weft length control with the help of robotic manipulation
In our implementation, we mount a camera on the top of fabric clamp to capture the position of the weft thread using the edge extraction based algorithm – see the window views at the upper-left corner of Fig.8(c) and (d). Two TM900 robotic arms are used to drive the slender hooks as their end effector, where the camera mounted on a robotic arm can help to successfully capture the weft thread by the hook. The robotic arms are placed in positions by checking the reachability map to ensure that the entire motion process avoids encountering the singularity poses. This robot-assisted tightening operation can help form stitches with better dimension control on the resultant fabrics. Note that using two robotic arms here is primarily for the sake of flexibility in implementation. The similar function can be realized by using either a single robotic arm or a well-designed sophisticated mechanism, which however could be restricted by the working envelope or the possible motion flexibility.
@@ 3.5 Control system
Figure 10: Given a target surface, we first generate a geodesic distance-field by using the central line of the piece as source (a). A stitch mesh for short-row shaping based knitting as shown in (b) can be generated from the isocurves of the geodesic distance-field by optimized quadrangles / triangles (details can be found in [21]). The stitch mesh can directly map into (c) a knitting map $K(\cdot,\cdot)$, which is converted into (d) a weaving map $W(\cdot,\cdot)$ and the machine code with its structure / syntax as explained in (e) to supervise the operations of our weaving machine.
@@ 4.1 Knitting map generation
The stitch mesh generated by the algorithm in [21] can satisfy all these requirements. As a result, every stitch can be mapped to a location in the 2D knitting map by its row index and its location in the row – see Fig.10(b, c) for an example.
@@ 4.2 Converting knitting map into weaving code
Figure 11: We tested our 3D surface weaving machine on four different examples, where their weaving maps are given in (a). The results of physical fabrication are shown in (b). Our weaving system enables the flow of electric current over the freeform surfaces of fabrics by using conductive threads in (c). The measurements obtained with the avometer represent the resistance between two ends of the weft thread, which indicates the continuity of the conductive weft thread.
@@
Trgl. Thread Num. Comp. Time (sec.)
@@
 Tests with different number of threads were conducted to demonstrate the scalability of our computational method although the maximally allowed number of warp threads on our prototype machine is only 100.
 The computing time of W-code generation includes both the step of weaving map generation and the step for converting it into W-code.
@@ 5.1 Statistics of computation and fabrication
The computational statistics of these examples have been listed in Table 1, where the examples are tested by using threads in two different resolutions. As it can be observed, the computation on all tests can be completed within 10 seconds and the major bottleneck is the step of knitting map generation [21]. The complexity of the knitting map generation algorithm is $O(AB\log B)$ with $A$ and $B$ being the maximal number of stitches along the row and the column directions on the knitting map respectively.
@@ 5.3 Analysis of geometric errors: 3D shape
One of the most important purposes of this work is to reach a high level of geometric accuracy on the 3D surface shape of the produced woven fabrics. In the conventional textile industry, a flattening algorithm (e.g., [25, 26]) was employed to compute a (trimmed) 2D panel shape to fit onto the desired 3D object (e.g., human bodies in the apparel industry). However, as 3D surfaces are in general not flattenable (or named as developable in geometric modeling), wrinkles will appear at regions where the curvature is large and leftovers are inevitable. For a better fit, smaller pieces of the flat-woven fabric are generated and stitched together, which involved time-consuming sewing processes that are labor-intensive (e.g., [27, 28, 29]).
We conducted experimental tests to study the shape approximation errors on a variety of models. Both the skinned shape by flat woven fabrics and the shape from 3D surface weaving are compared with the target shape (as the ground truth). A structured light based 3D scanner [30] is employed to generate the point clouds for each surface, which is also placed into a well-registered pose w.r.t. the ground truth shape. The shortest distances between the sampled points on the given surface to the scanned point clouds are employed as shape approximation errors, which are visualized as color maps and histograms in Fig.14. From the results of error analysis, we can easily conclude that models fabricated by our 3D surface weaving have much smaller errors in all examples – i.e., mainly less than 1mm as can be observed from the histograms. When the surface to be fabricated becomes more and more complex, the geometric error on fitting results by flat woven fabrics will increase significantly (see the triple-peak model in Fig.14 for an example).
@@ 5.6 Continuity of Threads
Our 3D surface weaving machine can help produce functional fabrics with many interesting applications in mechatronics. When conductive threads are used (as already demonstrated in Fig.11(c)), our 3D surface weaving approach enables the flow of electric current over the surfaces of fabrics so that they can work like soft PCBs. This has many potential applications such as functional garments [31], motion sensing [32], and ‘skin’ of cobot [33]. To enable this function, continuity on both the warp and the weft threads is required. While forming a 3D freeform shape by our surface weaving technique, we use a single thread for weft and multiple continuous threads for warp directions. Continuity is preserved on the resultant woven fabrics.
@@ 5.7 Limitations
One major limitation of our current implementation is that two robotic arms are employed for realizing the length control of weft threads. This significantly increases the cost and the complexity of the manufacturing system. A possible future work is to utilize sword bar mechanical structures to control the length and distribution of the weft thread while also using a camera for visual feedback and control.
@@
Figure 19: An example of vest-front fabric made by using 1) bare stainless steel wire for warp threads and 2) silicone insulated copper wire for weft threads (left), which can be developed into a touch sensing interface by capacitance [34]. (Middle) the distribution of touch that is sensed by the grid of conductive threads, and (right) the corresponding touch visualized on the 3D surface.
In the examples listed above, the threads with high stiffness (such as optical fibres) are only employed as warp threads. This is because the weft thread needs to form U-turn at the end of each row. However, we can employ the in-lay technique to ‘insert’ an additional thread (with high bending-stiffness) between those neighboring rows formed by a continuous weft thread. An example can be found in Fig.18. Note that the in-lay fabrication needs manual operations on our prototype machine in our current experiment. However, when the U-turn at the end of each row is tolerable on materials with high bending-stiffness (e.g., metal wires), our method can produce fabrics by using these stiff materials in both warp and weft threads – see Fig.19 for an example by using these conductive grids to form a capacitance-based touch sensing interface [34]. Lastly, the abrasion of warp threads caused by the reed’s motion requires the threads to be waxed to prevent damage to its ‘core’.
Figure 20: An example of failure case. (a) Local maximum will be generated when selecting the middle curves as source of the the geodesic distance field. On a field with local maximum located inside the surface patch, closed loops of isocurves will be formed in which case our algorithm of weaving map generation will fail. (b) Weaving map generation algorithm can only be applied after adding darts from the local maximum to place it on the boundary of a surface patch. This is how the Triple-Peak example shown in Figs.11 and 14 was fabricated.
There are also some limitation on the software aspect. Our current method of weaving-map generation heavily relies on selecting a good curve as source to generate geodesic distance-field. Different maps including failure cases can be generated when using different sources. One failure scenario occurs when the local minimum or maximum appears after applying the geodesic distance field, leading to the generation of isocurves in closed loops for weft threads (see Fig.20(a) for an example). To resolve this issue, we need to cut the given model by adding darts so that pushing the local minimum or maximum to the boundary of a surface patch to successfully generate the weaving map (as shown in Fig.20(b)). One possible future research is to automate the source selection while preventing the failure cases and reducing the distortion of stitch meshes constructed from the geodesic distance-field.